OpenStudy (anonymous):

Rachel went hiking for one hundred fifty minutes. Express this time in hour and minutes. A)1 hour 5 minutes B)1 hour fifty minutes C)2 hours 5 minutes D)2 hours thirty minutes

OpenStudy (azureilai):

60 minutes in an hour, so 150/60 = 2 2*60=120 150-120=30 minutes 2 hours and 30 minutes